One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by psbhowmick(878)

Let the first number = x.
Then the other number = 12 more than 'x' = (x+12).
Clearly, 'x' is the smaller number and (x+12) is the larger.
So the sum of the smaller number and twice the larger number = x + 2(x + 12) = 39
i.e. x + 2(x + 12) = 39
x + 2x + 2*12 = 39
3x + 24 = 39
3x = 39 - 24
3x = 15
x = 15/3
x = 5
So the smaller number = x = 5 and the larger number = (x + 12) = 5 + 12 = 17

Answer: 17